# Lost in Sequence: Do Large Language Models Understand Sequential Recommendation?

### v3

@@ 2.1.2. LLM for Sequential Recommendation
1) Generative Approach (Next Item Title Generation). LLM4Rec models designed for Next Item Title Generation perform recommendations using instruction-based prompts as shown in Table 1. For a user , the candidate item set of user is represented as , where is a negative item set for user , and is the number of negative items. Based on the item interaction sequence of user , i.e., , and the candidate item set, i.e., , we write the input prompt following the format shown in Table 1. Note that we introduce two projection layers, i.e., and , each of which is used to project item embeddings and user representations extracted from a pre-trained CF-SRec into LLMs, respectively. Following the completed prompts shown in Table 1, LLMs are trained for the sequential recommendation task through the Next Item Title Generation approach. Note that TALLRec, LLaRA, and CoLLM use LoRA (Hu et al., 2022) to finetune LLMs aiming at learning the sequential recommendation task, while A-LLMRec only trains and without finetuning the LLMs with LoRA. Please refer to the Appendix B.1. for more details on the projection layers as well as prompt construction.
2) Retrieval Approach (Next Item Retrieval). As shown in Table 2, we use and to denote prompts for users and items, respectively. Unlike the Next Item Title Generation approach where LLMs directly generate the title of the recommended item, the Next Item Retrieval approach generates item recommendations by computing the recommendation scores between user representations and item embeddings. More precisely, it introduces learnable tokens, i.e., [UserOut] and [ItemOut], to aggregate information from user interaction sequences and items, respectively. The last hidden states associated with the [UserOut] and [ItemOut] are used as user representations and item embeddings, denoted and , respectively, where denotes the token embedding dimension of LLM. Please refer to Appendix B.2. for more details on how the user representations and item embeddings are extracted as well as the prompt construction for compared models.
@@ 2.2. Evaluation Protocol
In our experiments on LLMs’ sequence comprehension, we employed the leave-last-out evaluation method (i.e., next item recommendation task) (Kang and McAuley, 2018; Sun et al., 2019; Tang and Wang, 2018a). For each user, we reserved the last item in their behavior sequence as the test data, used the second-to-last item as the validation set, and utilized the remaining items for training. The candidate item set (i.e., test set) for each user in the title generation task is generated by randomly selecting 19 non-interacted items along with 1 positive item following existing studies (Zhang et al., 2023; Kim et al., 2024). Similarly, for the next item retrieval task, we randomly select 99 non-interacted items along with 1 positive item as the candidate item set (i.e., test set) for each user.
@@ 2.3.1. Shuffled Training
Table 3 and Table 4 show the performance of various models when adapted to the Next Item Retrieval approach and the Next Item Title Generation approach, respectively. We have the following observations: 1) CF-SRec (i.e., SASRec), suffers from substantial performance degradation when the training sequences are shuffled as expected, whereas LLM4Rec models generally exhibit minimal changes in performance. This indicates that LLMs struggle to leverage sequential information, as eliminating original sequential dependencies through shuffling does not significantly impact their performance. 2) Some LLM4Rec models even show improved results despite being trained with shuffled sequences. We conjecture that in some cases random shuffling of interaction sequences during training can introduce short-term co-occurrence patterns that may coincidentally lead to improved performance. Combined with the fact that LLMs struggle to capture long-term dependencies (Liu and Abbeel, 2024), we argue that LLM4Rec models struggle to capture the sequential dependencies within the interaction sequence. 3) To investigate further on the behavior of models, we plot a test performance curve for each model during training (Figure 7 in Appendix F.1). Our expectation is that a model that understands the sequence would train more rapidly when trained with the original sequences, as the sequential information inherent in the item interaction sequence is helpful for recommendation. In other words, we expect the ratio of the slopes of the test performance curve (i.e., original vs. shuffled) to be larger for a model that understands the sequence. We observe from Figure 7 that while the ratio of SASRec is relatively high, that of LLM4Rec models are close to 1, indicating that LLM4Rec models fail to fully leverage the sequential information inherent in users’ item interaction sequences.
@@ 3.1. Distilling Sequential Information
where MSE is the mean squared error loss and both and are 2-layer MLPs , i.e., and , that are trainable. This simple distillation framework enables LLMs to generate user representations that effectively capture and reflect the sequential information inherent in users’ item interaction sequences. The prompt used in LLM-SRec is provided in Table 2, where user representations are derived from LLMs based on interacted item titles and collaborative information. In Appendix F.2, we empirically compare our prompt with another prompt that explicitly incorporates user representations, i.e., CoLLM and A-LLMRec, demonstrating the effectiveness of LLM-SRec despite the absence of user representation in the prompt. Additionally, Appendix F.3, we conducted experiments using a contrastive learning approach as an objective for distillation instead of the MSE loss.
@@ 3.2. Preventing Over-smoothing
The uniformity loss ensures that user representations of different users are uniformly distributed across the normalized feature space on the hypersphere, preserving both separation and informativeness. In Appendix F.4, we measure the pairwise distances between all user representations to validate the effectiveness of in mitigating the over-smoothing problem.
@@
It is important to note that LLM-SRec does not require any additional training for either the pre-trained CF-SRec or LLMs during its training process. Instead, LLM-SRec only optimizes a small set of MLP layers (i.e., , , and ) and two LLM tokens (i.e., [UserOut] and [ItemOut]). Therefore, LLM-SRec achieves faster training and inference time compared to existing LLM4Rec baselines, including TALLRec, LLaRA, CoLLM, and A-LLMRec, results are described in Sec. 4.3.1. Furthermore, for training efficiency, we only consider the last item in for each user to minimize Equation 4 (Kim et al., 2024). That is, for each user , we predict the last item given the sequence . In Appendix F.5, we also show the results when considering all items, i.e., auto-regressive learning.
@@ 4. Experiments
Datasets. We conduct experiments on four Amazon datasets (Hou et al., 2024a), i.e., Movies, Scientific, Electronics, and CDs. Following prior studies (Kang and McAuley, 2018; Sun et al., 2019), we use five-core datasets consisting of users and items with a minimum of five interactions each. The statistics for each dataset after preprocessing are summarized in Table 10 in Appendix C.
Baselines. We compare three groups of models as our baselines: models that use only interaction sequences (CF-SRec: GRU4Rec (Hidasi et al., 2015), BERT4Rec (Sun et al., 2019), NextItNet (Yuan et al., 2019), SASRec (Kang and McAuley, 2018)), Language Model based models (LM-based: CTRL (Li et al., 2023a), RECFORMER (Li et al., 2023c)), and Large Language Model based models (LLM4Rec: TALLRec (Bao et al., 2023), LLaRA (Liao et al., 2024), CoLLM (Zhang et al., 2023), A-LLMRec (Kim et al., 2024)). For fair comparisons, we implemented all LLM4Rec baselines with Next Item Retrieval approach. Details regarding the baselines are provided in Appendix D.
@@
Implementation Details. For fair comparisons, we adopt pre-trained LLaMA 3.2 (3B-Instruct) as the backbone LLM for all LLM4Rec baselines (i.e., TALLRec, CoLLM, LLaRA, and A-LLMRec) including LLM-SRec. Similarly, SASRec serves as the pre-trained CF-SRec for CoLLM, LLaRA, A-LLMRec, and LLM-SRec. Please refer to the Appendix E for more details regarding the hyper-parameters and train settings.
@@ 4.1.1. Overall performance
| SASRec | TALLRec | LLaRA | CoLLM | A-LLMRec | LLM-SRec | |
| Electronics | 0.1002 | 0.1214 | 0.1225 | 0.1232 | 0.1262 | 0.1310 |
| Scientific | ||||||
| Electronics | 0.0974 | 0.1132 | 0.1174 | 0.1152 | 0.1217 | 0.1369 |
| CDs |
@@ 4.2. Ablation Studies
In this section, we evaluate the contribution of each component LLM-SRec. To analyze not only the contribution of each component in terms of the final performance but also its impact on the sequence understanding ability, we conduct experiments under the setting described in Sec. 2.3.1. In other words, we compare the performance of training with the original sequences versus training with shuffled sequences. Table 7 presents the following observations: 1) When both and are present (i.e., vanilla LLM-SRec in row (c)), the model consistently achieves the highest performance on the original sequence due to the benefits of the sequence understanding ability. Furthermore, compared with its variants (i.e., row (c) vs (a,b)), the performance of the vanilla LLM-SRec drops the most rapidly when shuffling is applied, ensuring that the vanilla LLM-SRec indeed comprehends and effectively utilizes sequential information. 2) In the absence of both and (i.e., row (a)), where the sequential knowledge extracted from CF-SRec is not distilled to the LLMs, the model exhibits the lowest performance on the original sequence. Additionally, even when random shuffling is applied, the performance drop is minor. This result suggests that the model lacks sequence understanding capability without and . 3) When only is removed, the model suffers from the over-smoothing problem as discussed in Sec. 3.2, leading to lower performance compared to LLM-SRec. However, since is still present, we observe a significant performance drop when applying random shuffling. This once more indicates that the model is endowed with the sequence understanding ability with .
@@ 4.3.2. Size of LLMs
In this section, we conduct a case study on the Electronics dataset to qualitatively validate the benefit of sequential knowledge and LLMs’ textual understanding. Figure 6 shows three cases highlighting the effect of sequential knowledge and textual knowledge on recommendation. The cases are categorized as follows: (a) only LLM-SRec provides correct recommendations, (b) LLM4Rec models (i.e., A-LLMRec and LLM-SRec) provide correct recommendations while SASRec fails, and (c) LLM-SRec and SASRec provide correct recommendations while an LLM4Rec (i.e., A-LLMRec) baseline fails. We have the following observations: 1) In case (a), the user’s preference shifts from cable-related items to products from ”Apple” brand. LLM-SRec correctly recommends ”Apple Pencil,” while SASRec captures the changing preference but fails to recognize the textual information ”Apple,” leading to a wrong recommendation of an ”Amazon Fire 7 Tablet.” On the other hand, A-LLMRec, which struggles to capture sequential information, recommends ”Audio Cable” based on textual knowledge of ”Cable” and ”Speaker.” This emphasizes the importance of leveraging both sequential and textual information. 2) In case (b), the user focused on ”BOSE” brand products. Both LLM-SRec and A-LLMRec, leveraging textual knowledge, successfully recommended the ”BOSE” earbuds, while SASRec, lacking textual information, recommended the ”SAMSUNG Galaxy Buds.” This highlights the importance of textual knowledge in generating accurate recommendations. 3) In case (c), the user’s preference shifts from ”Hosa Tech” cables to security camera. LLM-SRec and SASRec capture this preference shift and provide relevant recommendations, while A-LLMRec recommends another ”Hosa Tech” cable ignoring the sequential patterns. Those cases demonstrate that both textual and sequential information are crucial for accurate recommendations, showcasing the superiority of LLM-SRec in integrating both for improved performance. We have provided an additional case study in Appendix F.6.
@@ 5. Related Work
Sequential Recommender Systems. Recommendation systems primarily focus on capturing collaborative filtering (CF) to identify similar items/users. Matrix Factorization-based approaches, achieved notable success by constructing CF knowledge in a latent space (Mnih and Salakhutdinov, 2007; Chaney et al., 2015; He et al., 2017). However, in conjunction with CF knowledge, understanding dynamic evolution in temporal user preferences has become a powerful tool, leading to the development of collaborative filtering-based sequential recommenders (CF-SRec) (Kang and McAuley, 2018; Sun et al., 2019; Wu et al., 2019; Kim et al., 2023; Hidasi et al., 2015; Oh et al., 2023). Initial approaches combined Matrix Factorization with Markov Chains to model temporal dynamics (Rendle et al., 2010). Subsequently, neural network-based methods advanced sequential recommender systems, with GRU4Rec (Hidasi et al., 2015) leveraging recurrent architectures, while methods such as Caser (Tang and Wang, 2018b) and NextItNet (Yuan et al., 2019) adopted Convolutional Neural Networks (Krizhevsky et al., 2012). More recently, models such as SASRec (Yuan et al., 2019) and BERT4Rec (Sun et al., 2019), based on attention mechanisms, have demonstrated superior performance by focusing on the more relevant interaction sequences. These advancements underscore the importance of effectively modeling user behavior dynamics for improved recommendation accuracy.

### v4

@@ 2.1.2. LLM for Sequential Recommendation
1) Generative Approach (Next Item Title Generation). LLM4Rec models designed for Next Item Title Generation perform recommendations using instruction-based prompts as shown in Table 1. For a user , the candidate item set of user is represented as , where is a negative item set for user , and is the number of negative items. Based on the item interaction sequence of user , i.e., , and the candidate item set, i.e., , we write the input prompt following the format shown in Table 1. Note that we introduce two projection layers, i.e., and , each of which is used to project item embeddings and user representations extracted from a pre-trained CF-SRec into LLMs, respectively. Following the completed prompts shown in Table 1, LLMs are trained for the sequential recommendation task through the Next Item Title Generation approach. Note that TALLRec, LLaRA, and CoLLM use LoRA (Hu et al., 2022) to finetune LLMs aiming at learning the sequential recommendation task, while A-LLMRec only trains and without finetuning the LLMs with LoRA. Please refer to the Appendix A.1. for more details on the projection layers as well as prompt construction.
2) Retrieval Approach (Next Item Retrieval). As shown in Table 2, we use and to denote prompts for users and items, respectively. Unlike the Next Item Title Generation approach where LLMs directly generate the title of the recommended item, the Next Item Retrieval approach generates item recommendations by computing the recommendation scores between user representations and item embeddings. More precisely, it introduces learnable tokens, i.e., [UserOut] and [ItemOut], to aggregate information from user interaction sequences and items, respectively. The last hidden states associated with the [UserOut] and [ItemOut] are used as user representations and item embeddings, denoted and , respectively, where denotes the token embedding dimension of LLM. Please refer to Appendix A.2. for more details on how the user representations and item embeddings are extracted as well as the prompt construction for compared models.
@@ 2.2. Evaluation Protocol
In our experiments on LLMs’ sequence comprehension, we employed the leave-last-out evaluation method (i.e., next item recommendation task) (Kang and McAuley, 2018; Sun et al., 2019; Tang and Wang, 2018a). For each user, we reserved the last item in their behavior sequence as the test data, used the second-to-last item as the validation set, and utilized the remaining items for training. The candidate item set (i.e., test set) for each user in the title generation task is generated by randomly selecting 19 non-interacted items along with 1 positive item following existing studies (Zhang et al., 2023; Kim et al., 2024). Similarly, for the next item retrieval task, following common strategies (Sun et al., 2019; Kang and McAuley, 2018), we randomly select 99 non-interacted items along with 1 positive item as the candidate item set (i.e., test set) for each user.
@@ 2.3.1. Shuffled Training
Table 3 and Table 4 show the performance of various models when adapted to the Next Item Retrieval approach and the Next Item Title Generation approach, respectively. We have the following observations: 1) CF-SRec (i.e., SASRec), suffers from substantial performance degradation when the training sequences are shuffled as expected, whereas LLM4Rec models generally exhibit minimal changes in performance. This indicates that LLMs struggle to leverage sequential information, as eliminating original sequential dependencies through shuffling does not significantly impact their performance. 2) Some LLM4Rec models even show improved results despite being trained with shuffled sequences. We conjecture that in some cases random shuffling of interaction sequences during training can introduce short-term co-occurrence patterns that may coincidentally lead to improved performance. Combined with the fact that LLMs struggle to capture long-term dependencies (Liu and Abbeel, 2024), we argue that LLM4Rec models struggle to capture the sequential dependencies within the interaction sequence.
@@ 3.1. Distilling Sequential Information
where MSE is the mean squared error loss and both and are 2-layer MLPs , i.e., and , that are trainable. This simple distillation framework enables LLMs to generate user representations that effectively capture and reflect the sequential information inherent in users’ item interaction sequences. The prompt used in LLM-SRec is provided in Table 2, where user representations are derived from LLMs based on interacted item titles and collaborative information. In Appendix E.1, we empirically compare our prompt with another prompt that explicitly incorporates user representations, i.e., CoLLM and A-LLMRec, demonstrating the effectiveness of LLM-SRec despite the absence of user representation in the prompt. Additionally, Appendix E.2, we conducted experiments using a contrastive learning approach as an objective for distillation instead of the MSE loss.
@@ 3.2. Preventing Over-smoothing
The uniformity loss ensures that user representations of different users are uniformly distributed across the normalized feature space on the hypersphere, preserving both separation and informativeness.
@@
It is important to note that LLM-SRec does not require any additional training for either the pre-trained CF-SRec or LLMs during its training process. Instead, LLM-SRec only optimizes a small set of MLP layers (i.e., , , and ) and two LLM tokens (i.e., [UserOut] and [ItemOut]). Therefore, LLM-SRec achieves faster training and inference time compared to existing LLM4Rec baselines, including TALLRec, LLaRA, CoLLM, and A-LLMRec, results are described in Sec. 4.3.1. Furthermore, for training efficiency, we only consider the last item in for each user to minimize Equation 4 (Kim et al., 2024). That is, for each user , we predict the last item given the sequence . In Appendix E.3, we also show the results when considering all items, i.e., auto-regressive learning.
@@ 4. Experiments
Datasets. We conduct experiments on four Amazon datasets (Hou et al., 2024a), i.e., Movies, Scientific, Electronics, and CDs. Following prior studies (Kang and McAuley, 2018; Sun et al., 2019), we use five-core datasets consisting of users and items with a minimum of five interactions each. The statistics for each dataset after preprocessing are summarized in Table 10 in Appendix B.
Baselines. We compare three groups of models as our baselines: models that use only interaction sequences (CF-SRec: GRU4Rec (Hidasi et al., 2015), BERT4Rec (Sun et al., 2019), NextItNet (Yuan et al., 2019), SASRec (Kang and McAuley, 2018)), Language Model based models (LM-based: CTRL (Li et al., 2023a), RECFORMER (Li et al., 2023c)), and Large Language Model based models (LLM4Rec: TALLRec (Bao et al., 2023), LLaRA (Liao et al., 2024), CoLLM (Zhang et al., 2023), A-LLMRec (Kim et al., 2024)). For fair comparisons, we implemented all LLM4Rec baselines with Next Item Retrieval approach. Details regarding the baselines are provided in Appendix C.
@@
Implementation Details. For fair comparisons, we adopt pre-trained LLaMA 3.2 (3B-Instruct) as the backbone LLM for all LLM4Rec baselines (i.e., TALLRec, CoLLM, LLaRA, and A-LLMRec) including LLM-SRec. Similarly, SASRec serves as the pre-trained CF-SRec for CoLLM, LLaRA, A-LLMRec, and LLM-SRec. Please refer to the Appendix D for more details regarding the hyper-parameters and train settings.
@@ 4.1.1. Overall performance
| SASRec | TALLRec | LLaRA | CoLLM | A-LLMRec | LLM-SRec | |
| Electronics | 0.1002 | 0.1214 | 0.1225 | 0.1232 | 0.1262 | 0.1310 |
| Scientific | ||||||
| Electronics | 0.0974 | 0.1132 | 0.1174 | 0.1152 | 0.1217 | 0.1369 |
| CDs |
@@ 4.2. Ablation Studies
In this section, we evaluate the contribution of each component LLM-SRec. To analyze not only the contribution of each component in terms of the final performance but also its impact on the sequence understanding ability, we conduct experiments under the setting described in Sec. 2.3.1. In other words, we compare the performance of training with the original sequences versus training with shuffled sequences. Table 8 presents the following observations: 1) When both and are present (i.e., vanilla LLM-SRec in row (c)), the model consistently achieves the highest performance on the original sequence due to the benefits of sequence understanding ability. Furthermore, compared with its variants (i.e., row (c) vs (a,b)), the performance of vanilla LLM-SRec drops the most rapidly when shuffling is applied, ensuring that vanilla LLM-SRec indeed comprehends and effectively utilizes sequential information. 2) In the absence of both and (i.e., row (a)), where the sequential knowledge extracted from CF-SRec is not distilled to the LLMs, the model exhibits the lowest performance on the original sequence. Additionally, even when random shuffling is applied, the performance drop is minor. This result suggests that the model lacks sequence understanding capability without and . 3) When only is removed, the model suffers from the over-smoothing problem as discussed in Sec. 3.2, leading to lower performance compared to LLM-SRec. However, since is still present, we observe a significant performance drop when applying random shuffling. This once more indicates that the model is endowed with the sequence understanding ability with .
@@ 4.3.2. Size of LLMs
In this section, we conduct a case study on the Electronics dataset to qualitatively validate the benefit of sequential knowledge and LLMs’ textual understanding. Figure 6 shows three cases highlighting the effect of sequential knowledge and textual knowledge on recommendation. The cases are categorized as follows: (a) only LLM-SRec provides correct recommendations, (b) LLM4Rec models (i.e., A-LLMRec and LLM-SRec) provide correct recommendations while SASRec fails, and (c) LLM-SRec and SASRec provide correct recommendations while an LLM4Rec (i.e., A-LLMRec) baseline fails. We have the following observations: 1) In case (a), the user’s preference shifts from cable-related items to products from ”Apple” brand. LLM-SRec correctly recommends ”Apple Pencil,” while SASRec captures the changing preference but fails to recognize the textual information ”Apple,” leading to a wrong recommendation of an ”Amazon Fire 7 Tablet.” On the other hand, A-LLMRec, which struggles to capture sequential information, recommends ”Audio Cable” based on textual knowledge of ”Cable” and ”Speaker.” This emphasizes the importance of leveraging both sequential and textual information. 2) In case (b), the user focused on ”BOSE” brand products. Both LLM-SRec and A-LLMRec, leveraging textual knowledge, successfully recommended the ”BOSE” earbuds, while SASRec, lacking textual information, recommended the ”SAMSUNG Galaxy Buds.” This highlights the importance of textual knowledge in generating accurate recommendations. 3) In case (c), the user’s preference shifts from ”Hosa Tech” cables to security camera. LLM-SRec and SASRec capture this preference shift and provide relevant recommendations, while A-LLMRec recommends another ”Hosa Tech” cable ignoring the sequential patterns. Those cases demonstrate that both textual and sequential information are crucial for accurate recommendations, showcasing the superiority of LLM-SRec in integrating both for improved performance.
@@ 5. Related Work
Sequential Recommender Systems. Recommendation systems primarily focus on capturing collaborative filtering (CF) to identify similar items/users. Matrix Factorization-based approaches, achieved notable success by constructing CF knowledge in a latent space (Mnih and Salakhutdinov, 2007; Chaney et al., 2015; He et al., 2017; Kim et al., 2025). However, in conjunction with CF knowledge, understanding dynamic evolution in temporal user preferences has become a powerful tool, leading to the development of collaborative filtering-based sequential recommenders (CF-SRec) (Kang and McAuley, 2018; Sun et al., 2019; Wu et al., 2019; Kim et al., 2023; Hidasi et al., 2015; Oh et al., 2023; Choi et al., 2025). Initial approaches combined Matrix Factorization with Markov Chains to model temporal dynamics (Rendle et al., 2010). Subsequently, neural network-based methods advanced sequential recommender systems, with GRU4Rec (Hidasi et al., 2015) leveraging recurrent architectures, while methods such as Caser (Tang and Wang, 2018b) and NextItNet (Yuan et al., 2019) adopted Convolutional Neural Networks (Krizhevsky et al., 2012). More recently, models such as SASRec (Yuan et al., 2019) and BERT4Rec (Sun et al., 2019), based on attention mechanisms, have demonstrated superior performance by focusing on the more relevant interaction sequences. These advancements underscore the importance of effectively modeling user behavior dynamics for improved recommendation accuracy.